Hi Can someone explain in detail the answer ?

public class Example
{
public static void main(String [] args)
{
int a = 0, b = 4;
  
do
{
a = a + 1;
b = b - 1;
System.out.println (a + " " + b);
}
while (b > a);
System.out.println ("ALL DONE");
}
}

answer

1 3
2 2
ALL DONE
Answer #1

step 1:- declare and initialize variable a=0 and b=4

step 2:- execute do while loop first time then execute statement a = a + 1; and b = b - 1; after executing both statement a will be 1 (means a=0+1=1) and b will be 3 (means b=4-1=3), after that execute System.out.println (a + " " + b); and print 1 3 and check while condition while (b>a) means while(3>1), so condition is true and loop execute again. so final value of a will be 1 and b will be 2 after first iteration.

step 3:-  execute do while loop second time then execute statement a = a + 1; and b = b - 1; after executing both statement a will be 2 (means a=1+1=1) and b will be 2 (means b=3-1=2), after that execute System.out.println (a + " " + b); and print 2 2 and check while condition while (b>a) means while(2>2), so condition is false and loop will be terminate and execute next statement System.out.println ("ALL DONE"); and print ALL DONE